OpenStudy (anonymous):

How would I simplify something like this: (12a^3-20a^2)/(16a^2+8a) Please include steps :)

OpenStudy (anonymous):

\[12a^3-20a^2/16a^2+8a\]

OpenStudy (anonymous):

If this is a limit calculation, I would divide everything by a^3. Else, I would first divide everything by 4 to obtain:\[\frac{3a^3 - 5a^2}{4a^2 + 2a}\]Then I would divide by a to get:\[\frac{3a^2 - 5a}{4a + 2}\]Now you can do polynomial division if you wish, personally I am not a really fan of it.

OpenStudy (anonymous):

4(3a^3-5a^2) / 4(3a^2+2a) 3a^3-5a^2 / 3a^2+2

OpenStudy (anonymous):

correct is: 4(3a^3-5a^2) / 4(4a^2+2a) 3a^3-5a^2 / 4a^2+2a

OpenStudy (anonymous):

Ah right, you can factor out stuff:\[\frac{a(3a - 5)}{2(2a + 1)}\]

OpenStudy (anonymous):

Thank you all :)

OpenStudy (anonymous):

No problem :-)